OpenStudy (anonymous):

Find the diameter of a circle with a circumference of 50 mm. Round your answer to the nearest tenth. 16.0 mm 26.4 mm 8.4 mm 33.5 mm

OpenStudy (mertsj):

What is the formula for the circumference of a circle?

OpenStudy (anonymous):

Use the formula to equate to 50mm and find dimaeter

OpenStudy (anonymous):

3.14 times what ?

OpenStudy (anonymous):

times 50

OpenStudy (anonymous):

?

OpenStudy (raden):

use this formula : C = pi * d with C=circumference, d=diameter, and pi=3.14

OpenStudy (anonymous):

so 3.14*50?

OpenStudy (raden):

Circumference is given, so pi * d = 50 3.14 * d = 50 now, solve for d

OpenStudy (anonymous):

157

OpenStudy (raden):

3.14 * d = 50 divided by 3.14 to both sides, gives us d = 50/3.14 = ....

OpenStudy (anonymous):

0.0628

OpenStudy (anonymous):

15.923 my bad

OpenStudy (raden):

yes, or it can be 16 :)